Find the electric field in a charge free region if it is pointing along the z-direction repeat this for a region having a charge density.


1
Expert's answer
2021-11-07T19:59:26-0500

"E=\frac{kQ}{r^3}{r}"

"r=r\hat{k}"

Charge density

"\rho=\frac{Q}{A}"

"Q=\rho A"

Now electric field

"E=\frac{k\rho A}{r^3}r\hat{k}"